The freezing point of pure camphor is 178.4°C, and its molal bpt contant, Kf, is 40°C. kg/mol. Find the freezing point of a solution containing 1.50 g of a compound of molar mass 125 in 35 g of camphor.


1
Expert's answer
2021-11-15T10:50:13-0500

∆Tf= Kf×m×i

178×125/35×40= 15.89

15.89×1.5= 23.84°C